OpenStudy (anonymous):

PLEASE help?!?! Can you describe the transformation of the blue line and give an equation of this graph?! Please.

OpenStudy (anonymous):

Show me the graph

OpenStudy (anonymous):

I posted the pic with the post :) the pic should be right under the message.

OpenStudy (anonymous):

It dos not show up can you send it again?

OpenStudy (anonymous):

does*

OpenStudy (anonymous):

Hopefully this works

OpenStudy (anonymous):

blue line to the red line or vice versa?

OpenStudy (anonymous):

Vice versa. It's asking for the transformation from the red to the blue line and an equation. Your help is very appreciated!

OpenStudy (anonymous):

Ok. the basic graph is for the equation \[\sqrt{x}\]

OpenStudy (anonymous):

from the red to the blue they moved left one and up one.

OpenStudy (anonymous):

What is the transformation from the red to blue? And the equation?

OpenStudy (anonymous):

And the equation? Sorry, didn't mean to ask the same question again as soon as you answered it.

OpenStudy (anonymous):

the equation for the blue line would be \[\sqrt{x-1}\]

OpenStudy (anonymous):

the red would be \[\sqrt{x}-1\]

OpenStudy (anonymous):

I just covered this in my Algebra II course.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No problem!